Deployment team

→ Europe/Zurich
EVO - GridPP Deployment team meeting

EVO - GridPP Deployment team meeting

Jeremy Coles
Description
- This is the weekly DTEAM meeting - The intention is to run the meeting in EVO: http://evo.caltech.edu/evoGate/. Join the meeting in the GridPP Community area - The phone bridge number is +41 22 76 71400,. Th phone bridge ID is 124322 with code: 4880.
Minutes
    • 11:00 → 11:20
      Experiment problems/issues 20m
      Review of weekly issues by experiment/VO - LHCb - CMS - ATLAS - Other Site status ********** - Accounting: IC-LeSC (trailing); LANCS-HEP - SAM: Average weekly success is down to 75% from 85%. Several sites in maintenance. Birmingham looks unstable. Sheffield and Liverpool both managed 100% for the last week... - SL ATLAS tests: Bristol and Edinburgh have clear problems.
    • 11:20 → 11:35
      ROC & WLCG update 15m
      ops meeting ******************** - CIC dashboard has been unreachable at times in last week - Since the last update of GOCDB some old nodes with monitoring off in the previous version are now monitored. It is now not possible to switch off the monitoring on nodes for a production site. Consequently site admins must delete all old nodes in the GOCDB which are not used anymore. These nodes must be removed from the information system in order not to be monitored by SAM. In case of differences between GOCDB and BDII (node is registered in the BDII and not in GOCBD , or a is not in the GOCDB but in the BDII) sites will receive COD tickets. - Response on PPS identified bugs not being fixed in recent release: "Regarding the two bugs raised in preproduction, that was actually a communication flaw in the procedure, which was revealed because of the high rate of releases. The flaw has been fixed, so the risk for this kind of glitch is from now on reduced." - There were three releases to PPS last week, mainly dealing with software needed for CCRC08, details available in https://twiki.cern.ch/twiki/bin/view/EGEE/PPSReleaseNotes - There was a discussion about transparent interventions and how they are notified. Noted that GOCDB has an "at risk" setting but it was not clear how this was defined. - There were three releases on gLite3.1 + one release on gLite3.0 last week. The released software mainly dealt with software needed for CCRC08. Details available in: http://glite.web.cern.ch/glite/packages/R3.1/updates.asp http://glite.web.cern.ch/glite/packages/R3.0/updates.asp - LHCb requested each of the T1 sites involved in the CCRC exercise to setup some informative twikis (and provide the link!) where a detailed description of the SRM-v2 implementation is given. AAM ******* - Main thing to note is that Debian was seen as the next main platform for porting efforts. Ticket status *************** See linked page
      more information
    • 11:35 → 11:45
      Quarterly reports 10m
      - item delayed from last week - review of issues arising - T2 activities for this quarter
    • 11:45 → 11:55
      Team review 10m
      - review of blogs - discussion on current team activities - what are the current priorities
    • 11:55 → 12:05
      Actions review 10m
    • 12:05 → 12:10
      AOB 5m
      - The RB and gLite-WMS situation was discussed by the PMB yesterday. D Colling mentioned that an SL4 version is coming soon and that should be our trigger to look at moving away from the LCG-RBs which are still looking unreliable and getting negative comments from users. - There is a CCRC'08 F2F meeting next week at CERN: http://indico.cern.ch/conferenceDisplay.py?confId=26922 - There is a GDB next Wednesday: http://indico.cern.ch/conferenceDisplay.py?confId=20226. Any things to raise? Meeting chat content: [11:02:26] Andrew Elwell https://twiki.cern.ch/twiki/bin/view/LCG/DpmAdminGuide [11:04:12] Pete Gronbech joined [11:05:28] Graeme Stewart joined [11:10:46] Frederic BROCHU joined [11:11:38] Frederic BROCHU left [11:11:58] Frederic BROCHU joined [11:12:30] Frederic BROCHU [11:12:37] Frederic BROCHU left [11:12:39] David Colling joined [11:13:01] David Colling Hallo all, apologies for being late had to reboot [11:13:17] Frederic BROCHU joined [11:13:37] Frederic BROCHU left [11:14:50] Graeme Stewart http://svr017.gla.scotgrid.ac.uk/factory/logs/2008-01-29/serv03.hep.phy.cam.ac.uk_2119_jobmanager-lcgcondor-atlas/SubmissionLog [11:15:45] Frederic BROCHU joined [11:16:51] Andrew Elwell ANY plans to install ATLAS software onto ECDF? [11:17:51] Graeme Stewart yes, but no joy so far... [11:18:56] Pete Gronbech Graeme: can you confirm if RALPPD are certified for ATLAS production? [11:19:03] Graeme Stewart yes [11:19:08] Pete Gronbech and JET? [11:19:17] Jens Jensen left [11:19:25] Jens Jensen joined [11:19:48] Jens Jensen Ah that's better - I can hear again [11:25:33] Alessandra Forti joined [11:45:18] Greig Cowan sorry, had to leave briefly [11:45:45] Raja Nandakumar joined [11:46:55] Greig Cowan 31344 [11:47:00] Greig Cowan GGUS number [11:49:44] Andrew Elwell I haven't closed the glasgow one (that came in about an hour ago...) [11:56:23] Jens Jensen The srm2 is described in thr GridPP wiki..? [12:03:41] Jens Jensen Ah, connected again... [12:04:00] Jens Jensen Networks are really &#*&$*ed up this morning [12:15:24] Andrew Elwell left [12:15:25] Greig Cowan left [12:15:28] Duncan Rand left [12:15:31] David Colling left [12:16:12] Graeme Stewart left [12:18:55] Jens Jensen left